Two balls are drawn in succession without replacement from an urn containing

2 green balls and 6 yellow balls. Let M be the random variable representing the

number of yellow balls.


1
Expert's answer
2021-05-17T02:19:03-0400

Total number of balls=8

Since Two balls are drawn at random and M denote the number of yellow balls.


So M cane be 0,1,2


Probability distribution of M is-


"P(M=0)= \dfrac{^6C_0\times ^2C_2}{^8C_2}=\dfrac{1}{28}=0.0357

\\[9pt]

P(M=1)= \dfrac{^6C_1\times ^2C_1}{^8C_2}=\dfrac{12}{28}=0.4287


\\[9pt]
P(M=2)= \dfrac{^6C_2\times ^2C_0}{^8C_2}=\dfrac{15}{28}=0.5357
\\[9pt]"